OpenStudy (anonymous):

y varies jointly with p and q. When p = 4 and q = 5, y = 60. Find y when p = 5 and q = 8.

OpenStudy (anonymous):

@Brittni0605 could you help me if your not busy? Thank You!

OpenStudy (anonymous):

Hmm, gimme a sec plz

OpenStudy (anonymous):

OKay thaks

OpenStudy (anonymous):

y varies jointly with p and q. the formula would be y=kpq 60=k(4)5 solve for k then Find y when p = 5 and q = 8.

OpenStudy (anonymous):

can you do the next step?

OpenStudy (anonymous):

@yazthespaz , you can follow @mark_o. I need a few min, sorry

OpenStudy (anonymous):

y= k(5)(8)= 40 so the answer is y = k (40)

OpenStudy (anonymous):

which will equal y = 40

OpenStudy (anonymous):

@Brittni0605 okay, its fine

OpenStudy (anonymous):

yes correct but did you do how to solve for k earlier

OpenStudy (anonymous):

.60=k(4)5

OpenStudy (anonymous):

yes thanks

OpenStudy (anonymous):

60=k(4)5 60=k(20) so 60/20=k k=__? ok good luck now it seems you got now the correct answer... :D